# Quality assurance tests and techniques to investigate and improve hermeticity of ATLAS Phase II Resistive Plate Chamber detectors

### v1

@@ Abstract
The Phase II upgrade of the ATLAS Muon Spectrometer will involve the installation of approximately 1000 next-generation Resistive Plate Chamber (RPC) singlets. This upgrade aims to enhance detector coverage, increase hit efficiency, and improve timing precision, ultimately strengthening the precision and robustness of the muon trigger system. The upgrade is essential for maintaining the performance of the muon spectrometer in the high-luminosity environment of the HL-LHC, where increased radiation and event rates pose significant challenges. Currently, detector production is underway, with gas gaps being commercially produced in Italy. To ensure the integrity of these gas gaps, especially their mechanical properties and gas tightness, several investigative techniques have been proposed and implemented.
@@ 1 Introduction
The new chambers are expected to significantly enhance the performance of ATLAS’s Muon system. By increasing the RPC-based muon trigger system from six to nine layers, the robustness of the system will improve, and the overall detector acceptance will rise from 78% to 92%, reaching up to 96% when combining Barrel Inner (BI) and Barrel Outer (BO) coincidences. The introduction of these new RPCs, featuring a thin gas gap (1 mm), will enable a timing resolution of approximately 0.4 ns for BI-RPCs and an improvement in legacy RPCs from 2 ns to 1.1 ns with upgraded readout electronics, thereby enhancing the accuracy of time-of-flight (ToF) measurements.
@@
Figure 3: Example of a thermal cycle performed in the Cambridge climate chamber over several hours, with recorded temperature conditions (top). Sensors at multiple positions – including within the gas gap inlet and outlet – track the temperature response in relation to the separately controlled climate chamber conditions. Recorded humidity conditions during the same thermal cycle (bottom), displaying relative humidity measurements at the gas gap inlet and outlet, as well as positions near the inlet and outlet on the outside of the gas gap. Additional sensors on the front and back surfaces provide a broader view of the climate room ambient conditions.
@@
The gas gaps used in these RPCs are primarily produced by General Tecnica Engineering (GTE) in Italy (with other potential production sites being considered in Germany and China) and are a core part of the performance and longevity of the detectors [5]. Each gas gap consists of a gas gap sandwiched between two High-Pressure Laminate (HPL) bakelite plates with a thin external graphite coating. After production, these gas gaps undergo several quality assurance tests, including surface resistivity measurements [6]. The study presented in this paper specifically focuses on evaluating the integrity and durability of these gas gaps under mechanical stress – a key aspect of assessing its reliability over a timespan of more than a decade during the operation of the HL-LHC.
@@
In this study, we describe our Quality Assurance (QA) tests for the gas gaps of Phase II RPC. The thermal cycling (TC) technique is central to these QA tests. This technique applies mechanical stress to the gas gaps by exploiting variations in the thermal expansion coefficients of the constituent materials of the gas gaps through systematic temperature variations. The key advantage of this technique is that it applies stresses in a controlled manner and that these stresses are focused on the interfaces between the different constituents of the gas gap like e.g., high-pressure laminate (bakelite) and glue, where mechanical failure is most likely to occur. Another advantage of this technique is that it can simulate the ageing of the gap in an environment with relatively stable environmental conditions over a decade-long timespan on a short timescale of a couple of days by using a sufficiently large window for the temperature excursions in the TC protocol. Similar TC procedures are used by ATLAS in the quality control and QA of the ITk strip modules [7] and serve as an inspiration for our tests. However, unlike for the ITk that looks for “trivial” mechanical failure modes like module cracking, our tests aim to reveal the mechanical resilience and potential failure points of the gas gaps in view of gas tightness, providing valuable insights into their structural integrity over time. The work presented here has supported improvements in gas gap design, thereby reducing emissions of RPC gases through leaks that could potentially develop over time during RPC operation and significantly contribute to global warming.
@@
Our QA tests through TC were carried within the climate-controlled chambers at both University of Cambridge and the INFN Frascati laboratories. These tests were conducted under conditions of down to $-33^{\circ}$ and up to $+35^{\circ}$ at Cambridge and down to $-20^{\circ}$ and up to $+20^{\circ}$ at Frascati. Between the TCs, the gas gaps were diligently monitored for their mechanical resilience and gas tightness, and regular searches for leak sites were carried out.
@@ 2 Methodology and Test Setup
Thermal cycling tests were conducted to assess the gas gaps under varying temperature conditions. As part of the procedure, gas gaps were initially evaluated for existing leaks by inflating them to an overpressure of either 3 or 5 mbar. The gas gaps are rated for up to 10 mbar overpressure, they were typically tested at 3 mbar. The use of 5 mbar in addition to 3 mbar was intended to expose small leaks that may not be detectable within a few minutes at a lower overpressure of 3 mbar, as will be clarified in subsequent discussions. After inflation, the pressure drop within the gas gaps was monitored using a differential pressure manometer – Kane 3200 Differential Pressure Manometer – over a three-minute period. Based on empirical data from a sample of $\mathcal{O}(100)$ gas gaps at CERN, acceptable values for pressure drop were set at $\Delta$P $<$ 0.01 mbar per minute, establishing a benchmark for gas tightness. Following this initial leak assessment, thermal cycling tests were performed, and additional leak rate measurements were taken between thermal cycles to evaluate any changes in the integrity of the gas gaps.
@@
Figure 2 shows the experimental setup within the climate chamber room at the University of Cambridge. This setup includes five ATLAS Phase II BIS-type gas gaps and one complete RPC singlet, all placed in a climate-controlled environment designed to facilitate TC. The chamber room, equipped with two compressors, allows temperature variations from $-40^{\circ}$ to $+35^{\circ}$, though for our tests, it was operated within a range of $-33^{\circ}$ to $+35^{\circ}$. An example of a thermal cycle between $-33^{\circ}$ and $+25^{\circ}$, conducted using this setup, is shown in Figure 3 (top). Here, it can be observed that all sensors closely track the climate room’s temperature variations. Meanwhile, Figure 3 (bottom) illustrates the relative humidity at various points on one of the gas gaps. The sensors labelled Inlet and Outlet are housed in 3D-printed fittings connected to gas lines, recording internal gas conditions within the gap. Notably, the inlet sensor shows lower humidity than the outlet, as dry CO2 gas was continuously flushed through the gas gap, and the outlet sensor captures a slight humidity increase due to exposure to the chamber’s humid environment. This small exchange of moisture with ambient air is further verified by the Near Inlet and Near Outlet sensors. These sensors start with similar humidity levels; however, the Near Outlet sensor shows a gradual decrease in humidity as it interacts with the gas flow, which becomes more humid as it picks up moisture within the gas gap. The Front and Back sensors provide additional readings of the external humidity surrounding the gas gap and within the broader climate room environment.
@@
Figure 4: The BIL-type gas gap undergoing thermal cycling tests at INFN Frascati (top). Recorded conditions within the test chamber (bottom), illustrating temperature variations (black) between +30C and -20C across multiple thermal cycles, along with dew point variations (red), indicating humidity changes within the test chamber.
@@ 3 Test Results
Thermal cycling tests were conducted on five gas gaps at Cambridge and eight gas gaps at INFN Frascati, with pressure drops recorded after every three to four cycles, followed by a detailed assessment at the end of the test campaign. The Barrel Inner Small (BIS) type gas gap samples, tested at Cambridge, involved a total of 16 cycles, executed in the following order with each cycle followed by a leak measurement:
@@
Table 1: Comparison of leak rates measured initially at CERN, and after the TC campaign at Cambridge.
Gas Gap Leak rate before TC Leak rate after TC
@@
The main finding of the QA tests carried out at Cambridge is that the TC protocol ranging from $-33^{\circ}$ and $+35^{\circ}$ C is found to have a notable ageing effect on the gas gaps, as shown in Table 1. The results after the TC protocol are also presented in Figure 5.
All gaps have developed a higher leak rate after the TC. Most notably, gas gap BIS2A-28/23, one of the two gas gaps that were already exhibiting leaks at CERN, has developed quite a sizeable leak. This is a clear warning sign that this gap would not age well if deployed for the HL-LHC, and hence should never be installed. On the other side, the gas gap BIS2A-08/23 that leaked initially, retained a relatively stable albeit notable leak rate. Finally, gas gap BIS2A-11/23, one of the three gaps that was not leaking initially, displayed some minor but below threshold leaks, while gaps BIS2A-01/23 and BIS2A-12/23 developed some leaks after TC. This requires additional investigation to pinpoint potential leakage sources. Initial measurements were taken over short intervals using a pressure manometer with a resolution of 0.01 mbar pressure drop.
For subsequent assessments, slightly longer-duration measurements were acquired with a semi-automatic setup, allowing for more precise monitoring. However, these measurements only provide initial observations for the first few minutes after inflating the gas gaps to 5 mbar and should not be compared with the long-duration measurements presented later in this study. The semi-automatic setup used multiple cross-calibrated pressure and temperature sensors connected to a Raspberry Pi, enabling sensor control and data acquisition. The setup allowed better monitoring of gas gap conditions throughout the testing process.
Figure 5: Leak rates for different gas gaps measured at Cambridge using a semi-automatic setup following thermal cycles.
To put the TC tests into perspective, the Araldite glue, a two-component toughened methacrylate adhesive used in gas gap assembly, has a thermal expansion coefficient of 85–122 $\times$ 10-6/C over the -30 to +30 C range [11, 10]. By comparison, bakelite has a very low thermal expansion coefficient of 22 $\times$ 10-6/[12], leading to stress at glue-bakelite interfaces and increasing rate of ageing when exposed to temperature variations.
@@
To illustrate the behaviour of a “bad” gas gap, Figure 7 shows pressure drop measurements for gap BIS2A-08/23 as a function of time after 16 TCs with stable temperature and humidity levels throughout. This gas gap exhibited a noticeable leak that worsened after thermal cycling. It is interesting to note that the leakrate is fairly high in the first $\mathcal{O}(10)$ min, and stabilises somewhat afterwards. A potential explanation for this could be that the micropores that have opened are partially sealed by not fully polymerised linseed oil within the gap.
@@
Another interesting result is presented in Figure 8, which depicts the leak rate for gas gap BIS2A-01/23 as a function of time after 16 TCs. It is apparent that over a period of more than 12h ambient pressure (top) and a subtle temperature differences (bottom) affect the leak rate. Specifically, a rapid initial decline in pressure is observed as the gas gap’s pressure drops from around 5 mbar, levelling off closer to 3 mbar. Similar to BIS2A-08/23 shown in Figure 7, the leak rate is significantly higher in the first $\mathcal{O}(10)$ min and decreases somewhat after that, although it remains intolerably high. Additionally, a positive slope in the leak rate aligns with minor temperature increases in the bottom graph, suggesting some thermal sensitivity, although the trend stabilises briefly before rising again with temperature.
Figure 8: Pressure drop over time for gas gap BIS2A-01/23 after 16 thermal cycles, recorded using the same setup and parameters described in Figure 6. This figure highlights threshold effects, suggesting that leaks may occur at higher pressures (above 3 mbar), whereas the gas gap maintains adequate performance at lower pressures (approximately 3 mbar or less). Relative humidity variations are also included for reference in the figure at the top while the lower bottom one includes temperature variations over a longer period recorded by Front, Back, and Inlet sensors. The Front and Back sensors track external climate room conditions, while the Inlet sensor monitors internal temperature fluctuations within the gas gap. The slight increase in pressure drop measurements toward the end is attributed to an overall temperature rise both inside and outside the gas gap.
@@
The rapid pressure drop in the first $\mathcal{O}(10)$ min observed for gas gaps BIS2A-08/23 and BIS2A-01/23 may be attributed to a subtle effect: micropores that open up with the pressure increase are partially sealed again by not fully polymerised linseed oil or not fully cured glue within the gap. If confirmed, this would be a clear indication where the production process of the gas gaps should be improved.
An independent test campaign was undertaken at INFN Frascati aimed to ensure comprehensive testing and included a systematic thermal cycling process with the following procedure
@@ 4 Fluorescent Tracer Tests
Figure 9: (top) A gas gap being filled with the fluorescent tracer solution using a bottle, with the height of the bottle gradually raised from the bottom, following Archimedes’ principle. (bottom) Gas gaps filled with the fluorescent tracer solution illuminated with UV light (external lights switched OFF), illustrating the process of leak detection. The concentration and path of the tracer (in green) highlighted in these images indicate the presence of leaks or potential defects near the gas inlet/outlets and within the gas distribution manifold.
@@ 5 Summary and Outlook
The current work, therefore, marks another step in advancing the quality assurance and quality control process for RPC gas gaps and provides promising evidence of the utility of thermal cycling for revealing structural weaknesses. Some gas gaps that were previously tested leak-tight developed leaks after thermal cycling, while existing leaks in others worsened. These observations highlight the method’s sensitivity in identifying structural weaknesses, particularly under temperature variations near the material and adhesive operating thresholds. The study revealed that leaks originated near specific regions, prompting targeted design improvements to mitigate these vulnerabilities.

### v2

@@ Abstract
The Phase II upgrade of the ATLAS Muon Spectrometer will involve the installation of approximately 1000 next-generation single Resistive Plate Chamber (RPC) detectors, also referred to as singlets. This upgrade aims to enhance detector coverage, increase hit efficiency, and improve timing precision, ultimately strengthening the precision and robustness of the muon trigger system. The upgrade is essential for maintaining the performance of the muon spectrometer in the high-luminosity environment of the HL-LHC, where increased radiation and event rates pose significant challenges. Currently, detector production is underway, with gas gaps being produced in Italy. To ensure the integrity of these gas gaps, especially their mechanical properties and gas tightness, several investigative techniques have been proposed and implemented.
@@ 1 Introduction
The new chambers are expected to significantly enhance the performance of ATLAS Muon system. By increasing the RPC-based muon trigger system from six to nine layers, the robustness of the system will improve, and the overall detector acceptance will rise from 78% to 92%, reaching up to 96% when combining Barrel Inner (BI) and Barrel Outer (BO) coincidences. The introduction of these new RPCs, featuring a thin gas gap (1 mm), will enable a timing resolution of approximately 0.4 ns for BI-RPCs and an improvement in legacy RPCs from 2 ns to 1.1 ns with upgraded readout electronics, thereby enhancing the accuracy of time-of-flight (ToF) measurements.
@@
Figure 3: Example of a thermal cycle performed in the Cambridge climate chamber over several hours, with recorded temperature conditions (top). Sensors at multiple positions – including within the gas gap inlet and outlet – track the temperature response in relation to the separately controlled climate chamber conditions. Additional sensors on the front and back surfaces provide a broader view of the climate room ambient conditions. (bottom) Recorded relative humidity conditions during the same thermal cycle within the gas gap inlet and outlet, as well as positions near the inlet and outlet on the outside of the gas gap.
@@
The gas gaps used in these RPCs are primarily produced by General Tecnica Engineering (GTE) in Italy (with other potential production sites being considered in Germany and China) [5]. Each gas gap consists of a gas gap sandwiched between two High-Pressure Laminate (HPL) bakelite plates with a thin external graphite coating. After production, these gas gaps undergo several quality assurance tests, including surface resistivity measurements [6]. The study presented in this paper specifically focuses on evaluating the integrity and durability of these gas gaps under mechanical stress – a key aspect of assessing reliability over a timespan of more than a decade during the operation of the HL-LHC.
@@
In this study, we describe our Quality Assurance (QA) tests for the gas gaps of Phase II RPC. The thermal cycling (TC) [^1] A thermal cycle refers to one complete variation between a positive temperature and a negative temperature, followed by a return to the initial temperature. technique is central to these QA tests. This technique applies mechanical stress to the gas gaps by exploiting variations in the thermal expansion coefficients of the constituent materials of the gas gaps through systematic temperature variations. The key advantage of this technique is that it applies stresses in a controlled manner and that these stresses are focused on the interfaces between the different constituents of the gas gap like e.g., high-pressure laminate (bakelite) and glue, where mechanical failure is most likely to occur. Another advantage of this technique is that it can simulate the ageing of the gap in an environment with relatively stable conditions over a decade-long timespan on a short timescale of a couple of days by using a sufficiently large window for the temperature excursions in the TC protocol. Similar TC procedures are used by ATLAS in the quality control and QA of the Inner TracKer (ITk) strip modules [7] and serve as an inspiration for our tests. However, unlike for the ITk that looks for mechanical failure modes like module cracking, our tests aim to reveal the mechanical resilience and potential failure points of the gas gaps in view of gas tightness, providing valuable insights into their structural integrity over time. The work presented here has supported improvements in gas gap design, thereby reducing emissions of RPC gases through leaks that could potentially develop over time during RPC operation and contribute to global warming.
@@
Our QA tests through TC were carried within the climate-controlled chambers at both University of Cambridge and the INFN Frascati laboratories. These tests were conducted under conditions of down to $-33^{\circ}$ and up to $+35^{\circ}$ at Cambridge and down to $-20^{\circ}$ and up to $+20^{\circ}$ at Frascati. Between the TCs, the gas gaps were monitored for their mechanical resilience and gas tightness, and tests were conducted to identify potential leak points.
@@ 2 Methodology and Test Setup
Thermal cycling tests were conducted to assess the gas gaps under varying temperature conditions. As part of the procedure, gas gaps were initially evaluated for existing leaks by inflating them to an overpressure of either 3 or 5 mbar. The gas gaps are rated for up to 10 mbar overpressure, they were typically tested at 3 mbar. The use of 5 mbar in addition to 3 mbar was intended to expose small leaks that may not be detectable within a few minutes at a lower overpressure of 3 mbar, as will be clarified in subsequent discussions. After inflation, the pressure drop within the gas gaps was monitored using Kane 3200 Differential Pressure Manometer over a three-minute period. Based on empirical data from a sample of gas gaps at CERN, acceptable values for pressure drop were set at $\Delta$P $<$ 0.01 mbar per minute, establishing a benchmark for gas tightness. Following this initial leak assessment, thermal cycling tests were performed, and additional leak rate measurements were taken between thermal cycles to evaluate any changes in the integrity of the gas gaps.
@@
Figure 2 shows the experimental setup within the climate chamber room at the University of Cambridge. This setup includes five ATLAS Phase II Barrel Inner Small (BIS)-type gas gaps and one complete RPC detector, all placed in a climate-controlled environment designed to facilitate TC. The chamber room, equipped with two compressors, allows temperature variations from $-40^{\circ}$ to $+35^{\circ}$, though for our tests, it was operated within a range of $-33^{\circ}$ to $+35^{\circ}$. An example of a thermal cycle between $-33^{\circ}$ and $+25^{\circ}$, conducted using this setup, is shown in Figure 3 (top). Here, it can be observed that all sensors closely track the climate room’s temperature variations. Meanwhile, Figure 3 (bottom) illustrates the relative humidity at various points on one of the gas gaps. The sensors labelled Inlet and Outlet are housed in 3D-printed fittings connected to gas lines, recording internal gas conditions within the gap. Notably, the inlet sensor shows lower humidity than the outlet, as dry CO2 gas was continuously flushed through the gas gap, and the outlet sensor captures a slight humidity increase due to exposure to the chamber’s humid environment. This small exchange of moisture with ambient air is further verified by the Near Inlet and Near Outlet sensors. These sensors start with similar humidity levels; however, the Near Outlet sensor shows a gradual decrease in humidity as it interacts with the gas flow, which becomes more humid as it picks up moisture within the gas gap. The Front and Back sensors provide additional readings of the external humidity surrounding the gas gap and within the broader climate room environment.
@@
Figure 4: BIL-type gas gap undergoing thermal cycling tests at INFN Frascati (top). Recorded conditions within the test chamber (bottom), illustrating temperature variations (black) between +30C and -20C across multiple thermal cycles, along with dew point variations (red), indicating humidity changes within the test chamber.
@@ 3 Test Results
Thermal cycling tests were conducted on five gas gaps at Cambridge and eight gas gaps at INFN Frascati, followed by a detailed assessment at the end of the test campaign. The BIS type gas gap samples, tested at Cambridge, involved a total of 16 cycles, executed in the following order with each cycle followed by a leak measurement:
@@
Table 1: Comparison of pressure drop rates measured initially at CERN, re-evaluated at Cambridge before thermal cycling (TC), and after the TC campaign at Cambridge. The measurements were performed using a differential pressure manometer.
Gas Gap Before TC After TC
@@
The main finding of the QA tests carried out at Cambridge is that the TC protocol is found to have a notable ageing effect on the gas gaps, as shown in Table 1. Since the measured leak rates in RPC gas gaps are directly related to pressure drop rates, the latter has been used as an alternative method for quantifying leak rates. As a result, these terms have been used interchangeably in the current work.
All gaps have developed a higher leak rate after the TC. Most notably, gas gap BIS2A-28/23, one of the two gas gaps that were already exhibiting leaks at CERN, has developed quite a sizeable leak. This is a clear warning sign that this gap would not age well if deployed for the HL-LHC, and hence should never be installed. On the other side, the gas gap BIS2A-08/23 that leaked initially, retained a relatively stable albeit notable leak rate. Finally, gas gap BIS2A-11/23, one of the three gaps that was not leaking initially, displayed some minor but below threshold leaks, while gaps BIS2A-01/23 and BIS2A-12/23 developed some leaks after TC. This requires additional investigation to pinpoint potential leakage sources. Initial measurements were taken over short intervals using a pressure manometer with a resolution of 0.01 mbar.
For subsequent assessments, slightly longer-duration measurements were acquired with a semi-automatic setup, allowing for more precise monitoring. The semi-automatic setup used multiple cross-calibrated pressure and temperature sensors connected to a Raspberry Pi, enabling sensor control and data acquisition. The setup allowed better monitoring of gas gap conditions throughout the testing process. The results after the TC protocol are presented in Figure 5. These measurements provide initial observations within the first few minutes after inflating the gas gaps to 5 mbar and are consistent with those obtained using the differential manometer (Table 1). However, these should not be compared with the long-duration measurements presented later in this study.
Figure 5: Pressure drop rates for different gas gaps measured at Cambridge using a semi-automatic setup following thermal cycles.
To put the TC tests into perspective, the Araldite glue, a two-component toughened methacrylate adhesive used in gas gap assembly, has a thermal expansion coefficient of 85–122 $\times$ 10-6/C over the -30 to +30 C range [9, 10, 11]. By comparison, bakelite has a very low thermal expansion coefficient of 22 $\times$ 10-6/[12], leading to stress at glue-bakelite interfaces and increasing rate of ageing when exposed to temperature variations.
@@
To illustrate the behaviour of a “bad” gas gap, Figure 7 shows pressure drop measurements for gap BIS2A-08/23 as a function of time after 16 TCs with stable temperature and humidity levels throughout. This gas gap exhibited a noticeable leak that worsened after thermal cycling. It is interesting to note that the leakrate is fairly high in the first $\mathcal{O}(10)$ minutes, and stabilises somewhat afterwards. A potential explanation for this could be that the micropores that have opened are partially sealed by not fully polymerised linseed oil within the gap.
@@
Another interesting result is presented in Figure 8, which depicts the leak rate for gas gap BIS2A-01/23 as a function of time after 16 TCs. It is apparent that over a period of more than 12h ambient pressure (top) and a subtle temperature differences (bottom) affect the leak rate. Specifically, a rapid initial decline in pressure is observed as the gas gap pressure drops from around 5 mbar, levelling off closer to 3 mbar. Similar to BIS2A-08/23 shown in Figure 7, the leak rate is significantly higher in the first $\mathcal{O}(10)$ minutes and decreases somewhat after that, although it remains intolerably high. Additionally, a positive slope in the leak rate aligns with minor temperature increases in the bottom graph, suggesting some thermal sensitivity, although the trend stabilises briefly before rising again with temperature.
Figure 8: Pressure drop over time for gas gap BIS2A-01/23 after 16 thermal cycles, recorded using the same setup and parameters described in Figure 6. This figure highlights threshold effects, suggesting that leaks may occur at higher overpressures (above 3 mbar), whereas the gas gap maintains adequate performance at lower overpressures (approximately 3 mbar or less). Relative humidity variations are also included for reference in the figure at the top while the lower bottom one includes temperature variations over a longer period recorded by Front, Back, and Inlet sensors. The Front and Back sensors track external climate room conditions, while the Inlet sensor monitors internal temperature fluctuations within the gas gap. The slight increase in overpressure measurements toward the end is attributed to an overall temperature rise both inside and outside the gas gap.
@@
The rapid pressure drop in the first $\mathcal{O}(10)$ minutes observed for gas gaps BIS2A-08/23 and BIS2A-01/23 may be attributed to a subtle effect: micropores that open up with the pressure increase are partially sealed again by not fully polymerised linseed oil or not fully cured glue within the gap. If confirmed, this would be a clear indication where the production process of the gas gaps should be improved.
An independent test campaign was undertaken at INFN Frascati to ensure comprehensive testing, including a systematic thermal cycling process with the following procedure
@@ 4 Fluorescent Tracer Tests
Figure 9: (top) A gas gap being filled with the fluorescent tracer solution using a bottle, with the height of the bottle gradually raised from the bottom. (bottom) Gas gaps filled with the fluorescent tracer solution illuminated with UV light (external lights switched OFF), illustrating the process of leak detection. The concentration and path of the tracer (in green) highlighted in these images indicate the presence of leaks or potential defects near the gas inlet/outlets and within the gas distribution manifold.
@@ 5 Summary and Outlook
The current work, therefore, marks another step in advancing the quality assurance and quality control process for RPC gas gaps, providing promising evidence of the effectiveness of thermal cycling in revealing structural weaknesses. Some gas gaps that initially tested leak-tight developed leaks after thermal cycling, while pre-existing leaks in others worsened. These observations highlight the role of thermal cycling in identifying structural vulnerabilities, particularly under temperature variations. The study also revealed that leaks originated near specific regions, prompting targeted design improvements to mitigate these weaknesses.